# Can Large Language Models be Effective Online Opinion Miners?

### v2

@@ Abstract
Ryang Heo    Yongsik Seo    Junseong Lee    Dongha Leethanks:   Corresponding author Yonsei University - DLI Lab {ryang1119, ysseo, brulee, donalee}@yonsei.ac.kr
@@ 1 Introduction
Despite these advances, existing approaches still face two critical limitations. (1) Underrepresentation of real-world input complexity: Previous benchmarks predominantly focus on single-sentence reviews Peng et al. (2019); Cai et al. (2021); Zhang et al. (2021) or preprocessed dialogue scenarios Li et al. (2023). However, in real-world online environments, user opinions appear in far more complex and structurally diverse formats. In practice, opinion streams span multi-party threaded discussions, long-form narratives with interleaved pros/cons, and domain-specific markers (e.g., emojis, slang, abbreviations) that introduce implicit sentiment signals (Figure 1 Lower). The absence of a setting that comprehensively captures these realistic and diverse forms of opinion expression makes it difficult to assess under what conditions and to what extent large language models (LLMs) can effectively perform opinion mining. This gap poses a significant challenge to evaluating the utility of LLMs and understanding their applicability to real-world applications.
@@ 4 Results and Discussion
In this section, we present the main findings of our study. Each subsection addresses the research question—Can LLMs serve as effective online opinion miners?—from various perspectives, supported by detailed experimental results and analyses.
@@ 4.1 RQ1: What makes it challenging for LLMs to extract structured opinions?
4.2 RQ2: How meaningfully can LLMs generate meaningful opinion insights?
@@ B.1 Evaluation Models
Open-sourced LLMs
@@ B.3 Opinion-centric insight generation
Although the A3CU metric assigns the review site example a high score (60.91) and the Reddit example a low score (10.12), our reference-free LLM-as-a-judge evaluation reveals that the model’s performance on the Reddit content is in fact stronger across several human-aligned dimensions—particularly Coverage and Specificity. This divergence highlights the limitations of purely reference-based, automatic metrics in capturing the nuanced, insight-oriented qualities of opinion insights. We therefore conclude that for OIG evaluation, combining automatic reference-based metrics with a reference-free, human-aligned judging protocol yields a more comprehensive and reliable assessment of LLMs’ true insight-generation capabilities.
@@ B.5 Prompts
  • Feature-centric opinion extraction: The prompt designed for Feature-centric opinion extraction is shown in Table 16.
  • Opinion-centric insight generation: The prompt designed for Opinion-centric insight generation is shown in Table 17.
  • Contextual Match: The prompt used for performing Contextual Match is provided in Table 18.
@@
Post #11518 (Author: Person 1): Y34 Infiniti M45.340hp V8 luxury sedan with strong, square styling. Roughly 9,000 imported over the two years on sale.Parts availability is a bitch. Comment #11519 (Author: Person 2, Reply to reply comment #11518): Good choice. An actual car that people dont talk about. There was also the M56 with a VK56 V8 that made 420 horsepower. Also rare. Heard one with an exhaust a while ago and it sounded pretty good. Comment #11520 (Author: Person 3, Reply to reply comment #11519): Yeah, the M56 is really cool and almost never gets talked about. Infiniti also continued the V8 when they renamed the M37M56 to the Q70, but a V8 Q70 is even rarer than an M56. Comment #11521 (Author: Person 4, Reply to reply comment #11518): And they look so rad. Quirky Japanese styling that wasnt super well received by the western market at the time, but few things look as good as a murdered out M45 Comment #11522 (Author: Person 5, Reply to reply comment #11518): Always loved these and thought of them as cars that should never have made it here stateside because of how JDM the styling is. Could be such a timeless car with some right modifications Comment #11523 (Author: Person 6, Reply to reply comment #11518): Most sinister looking car since the w109 Mercedes
@@
Took awhile but have grown from tolerating to appreciating this SUV. Replacement of a defective 12v battery and the dealer updating my car from the 2.0 to the 2.1 software resolved the initial issues that initially were bricking my XC60 B5. Second, I turned OFF the automatic software updates which you can do in check for software update submenu. Sticking with version 2.1 for now since it has made my car stable (at least on my B5 trim). 2.2 is available which includes Apple CarPlay but that version was JUST released and the initial bugs have not been worked out. So buggy that 2.2 is not available for an OTA update specifically for B5 trims since it causes known problems. While CarPlay is in high demand the native Google OS has been great and the Google map appears on the driver display which CarPlay currently wont. That said, recommend finding the software version that enables stability for your car/trim and then turn off automatic updates in the settings. For my XC60 B5 I am going to stick with 2.1 until forced to update or until the blogs indicate Volvo and Google are aligning better. Aside from the software gremlins this SUV is comfortable, super handsome in my r-design trim, and feels a touch unique.
@@
The Hyundai Kona has a very open and room-like interior, with some well-placed buttons. I’ve been on an ergonomics kick lately, closely scrutinizing the button placement and cockpit layouts of every car I get into. The 2024 Hyundai Kona has a uniquely open and homey-feeling interior, but one aspect of it that struck me was how perfect the heated seat and steering wheel buttons are. Before your eyes roll out of your head at the thought of praising something as mundane as a button, consider this: Heated seat and steering wheel controls are features you might use every day, multiple times a day, for months and months while you own a car. It’s fun to talk about max skidpad g’s and 0-60 times, but realistically, how often do you take your vehicle to its extremes? When we’re talking about what it’s like to live with a car day in and day out, ergonomic touches can make or break the enjoyability of ownership.
@@
Coverage: 4 Comprehensive Coverage of Main Issues: The insight summary effectively highlights the initial issues faced by the reviewer, namely the defective 12v battery and problematic software updates, and how these were resolved by upgrading to software version 2.1. These are the critical components of the review text. Inclusion of Recommendation and Software Concerns: The insight summary correctly encapsulates the reviewer’s advice against automatic updates due to the issues with version 2.2, emphasizing its buggy nature and incompatibility with B5 trims. It also mentions the preference to stay with version 2.1, another key opinion in the review. Minor Missing Details: Although the insight summary is generally accurate, it doesn’t mention the specific aspect of finding a stable software version and then turning off automatic updates, which adds a nuanced detail to the reviewer’s overall advice. This omission prevents the insight summary from achieving a perfect score.
@@
Specificity: 4 Resolution Details: The insight summary covers the resolution process involving the upgrade to software version 2.1 to address initial vehicle issues. It explicitly mentions the defects in the 12v battery and problematic updates, demonstrating concrete detail awareness about the user’s experience. Specific Features and Versions: The insight summary specifies the difference between software versions 2.1 and 2.2, noting the inclusion of Apple CarPlay in version 2.2. It captures the user’s preference for version 2.1 for stability reasons and references the native Google OS’s features, offering precise insights. General Impression and Recommendations: While largely specific in detail, certain generalizations are made, such as mentioning the comfort and uniqueness of the SUV without referencing specific features that contribute to these impressions. This is where the insight summary could benefit from slightly more precise detail, warranting a score of 4 rather than a 5.
@@
Insightfulness: 4 Comprehensive Coverage: The insight summary effectively covers the key points of the review, including the issue with the battery, the software problems, and the resolution through the 2.1 software update. It emphasizes the importance of stability and cautions against automatic updates. This communicates the core guidance intended by the reviewer. Value-Added Advice: By recommending a strategic approach to software updates and emphasizing the stability offered by version 2.1, the insight summary provides actionable insights that can directly inform decision-making for potential users of the XC60 B5, particularly those concerned about software reliability. Balanced Perspective: The insight summary maintains a balance by acknowledging the non-software-related positives of the vehicle, such as comfort and aesthetic appeal, which enriches understanding for readers interested in a broader evaluation of the SUV’s attributes. However, it could delve deeper into why the Google OS is particularly advantageous or why the SUV’s design stands out, which slightly limits its insightfulness further.
@@
Intent: 5 Tone and Intent Preservation: The insight summary effectively captures the reviewer’s tone of cautious appreciation towards the vehicle, highlighting the resolve to prioritize stability over newer, but potentially problematic, software updates. The advice against automatic updates and emphasis on waiting for software stability clearly mirrors the reviewer’s thoughtful, pragmatic approach. Emotional Nuances: The emotional essence of the original review, which oscillates between initial frustration and eventual appreciation, is well-preserved. The insight summary accurately reflects the transition from dissatisfaction due to battery and software issues to satisfaction with the SUV’s comfort and aesthetic appeal. Detail and Specificity: Key details concerning the software update issues, particularly the problems associated with version 2.2 and the decision to stick with version 2.1, are thoroughly covered in the insight summary. The review’s specific mentions of the native Google OS and Apple’s CarPlay availability are adequately noted, maintaining the original’s nuanced reflections.
@@
Faithfulness: 5 Accurate Representation: The insight summary accurately captures the main points and perspectives shared by the commenters without distorting their meaning. Each viewpoint, from aesthetics to practicality, is clearly represented as expressed in the original comments. No Omitted Key Details: The core opinions about the use of ICE hulls for EVs, the practical considerations for work vans, and the limited design flexibility due to utility needs are all present in the insight summary. This shows no significant omissions. Absence of Hallucinated Information: The insight summary does not introduce any information that was not present in the original text, ensuring that it stays true to the opinions and insights provided by the commenters. This maintains the integrity of the review discussion.
@@
Specificity: 5 Concrete Alignment with Comments: The insight summary effectively captures the conversation between different users by directly referencing the points they make. It notes Person 1’s dissatisfaction with ICE-based EV designs and preference for innovative startups, Person 2’s emphasis on practicality and part availability for work vans, Person 3’s mention of financial and practical factors, and Person 4’s blunt assertion about vans’ inherent boxiness. These details perfectly align with the provided comments, making the insight summary highly specific. No Overgeneralization: At no point does the insight summary veer into vague or generic statements. Each opinion is not only linked to the respective commenter but also articulated with clear explanation of their rationale, demonstrating a detailed and precise understanding of the user opinions. Point-wise Clarity: The insight summary maintains a structurally coherent format by presenting each perspective distinctly. The inclusion of each contributor’s viewpoint provides a comprehensive view of the entire discussion, acknowledging the nuances in opinions rather than making an overarching generalization. This results in a precise and informative overview.
Insightfulness: 4 Comprehensive Overview: The insight summary effectively synthesizes various perspectives from the commentary, covering not only the dissatisfaction with traditional designs from Person 1 but also contrasting practical views from others who value utility over aesthetics. This balanced presentation allows readers to see different angles of the ongoing design debate, providing a fuller context. Emphasis on Practical Factors: By highlighting the points about practicality—such as access to parts, financial logic, and the vans’ purpose—the insight summary goes beyond mere restatement of comments and offers a deeper understanding of the practical considerations influencing design decisions in the EV market. Concludes with Functional Insight: The insight summary concludes with Person 4’s point about the inherent boxy design of vans due to their function, driving home the idea that practicality often trumps aesthetic innovation in vehicle design. This insight allows readers to grasp why some design choices remain seemingly unchanged.
Intent: 5 Tone and Intent Preservation: The insight summary accurately captures the tone and intent of each comment within the discussion. Person 1’s dissatisfaction with the trend is clearly articulated, as is the practical-focused tone of Persons 2, 3, and 4. This indicates that the emotional elements of approval or disapproval were preserved effectively. Nuance and Specificity: The insight summary encapsulates the nuances in each participant’s perspective. It highlights Person 1’s preference for innovative design, contrasts this with the practical considerations highlighted by Person 2 and Person 3, and points out Person 4’s straightforward viewpoint regarding the inherent functionality of vans. Consistent Style and Balance: The insight summary maintains a balanced view and communicates both sides of the argument without straying from the original intent of the discourse. This illustrates a good level of detail while maintaining the integrity and style of the original review, indicating an effective preservation of stylistic and emotional essence.

### v3

@@ Abstract
Ryang Heo    Yongsik Seo    Junseong Lee    Dongha Leethanks:   Corresponding author Department of Artificial Intelligence, Yonsei University - DLI Lab {ryang1119, ysseo, brulee, donalee}@yonsei.ac.kr
@@ 1 Introduction
Despite these advances, existing approaches still face two critical limitations. (1) Underrepresentation of real-world input complexity: Previous benchmarks predominantly focus on single-sentence reviews Peng et al. (2019); Cai et al. (2021); Zhang et al. (2021) or preprocessed dialogue scenarios Li et al. (2023). However, in real-world online environments, user opinions appear in far more complex and structurally diverse formats. In practice, opinion streams span multi-party threaded discussions, long-form narratives with interleaved pros/cons, and domain‐specific markers (e.g., emojis, slang, abbreviations) that introduce implicit sentiment signals (Figure 1 Lower). The absence of a setting that comprehensively captures these realistic and diverse forms of opinion expression makes it difficult to assess under what conditions and to what extent large language models (LLMs) can effectively perform opinion mining. This gap poses a significant challenge to evaluating the utility of LLMs and understanding their applicability to real-world applications.
@@ 4 Results and Discussion
In this section, we present the main findings of our study. Each subsection addresses the research question “Can LLMs serve as effective online opinion miners?” from various perspectives, supported by detailed experimental results and analyses.
@@ 4.1 RQ1: What makes it challenging for LLMs to extract structured opinions?
4.2 RQ2: How effectively can LLMs generate opinion-centric insight?
@@ B.1 Evaluation Models
Open-source LLMs
@@ B.3 Opinion-centric insight generation
Although the A3CU metric assigns the review site example a high score (60.91) and the Reddit example a low score (10.12), our reference-free LLM-as-a-judge evaluation reveals that the model’s performance on the Reddit content is in fact stronger across several human-aligned dimensions—particularly Coverage and Specificity. This divergence highlights the limitations of purely reference-based, automatic metrics in capturing the nuanced, insight-oriented qualities of opinion insights. We therefore conclude that for OIG evaluation, combining automatic reference-based metrics with a reference-free, human-aligned judging protocol yields a more comprehensive and reliable assessment of LLMs’ true insight‐generation capabilities.
@@ B.5 Prompts
  • Feature-centric opinion extraction (FOE): The prompt designed for Feature-centric opinion extraction is shown in Table 16.
  • Opinion-centric insight generation (OIG): The prompt designed for Opinion-centric insight generation is shown in Table 17.
  • Contextual Match (CM): The prompt used for performing Contextual Match is provided in Table 18.
@@
Post #11518 (Author: Person 1): Y34 Infiniti M45.340hp V8 luxury sedan with strong, square styling. Roughly 9,000 imported over the two years on sale.Parts availability is a bitch. Comment #11519 (Author: Person 2, Reply to reply comment #11518): Good choice. An actual car that people dont talk about. There was also the M56 with a VK56 V8 that made 420 horsepower. Also rare. Heard one with an exhaust a while ago and it sounded pretty good. Comment #11520 (Author: Person 3, Reply to reply comment #11519): Yeah, the M56 is really cool and almost never gets talked about. Infiniti also continued the V8 when they renamed the M37M56 to the Q70, but a V8 Q70 is even rarer than an M56. Comment #11521 (Author: Person 4, Reply to reply comment #11518): And they look so rad. Quirky Japanese styling that wasnt super well received by the western market at the time, but few things look as good as a murdered out M45
@@
Took awhile but have grown from tolerating to appreciating this SUV. Replacement of a defective 12v battery and the dealer updating my car from the 2.0 to the 2.1 software resolved the initial issues that initially were bricking my XC60 B5. Second, I turned OFF the automatic software updates which you can do in check for software update submenu. Sticking with version 2.1 for now since it has made my car stable (at least on my B5 trim). 2.2 is available which includes Apple CarPlay but that version was JUST released and the initial bugs have not been worked out. So buggy that 2.2 is not available for an OTA update specifically for B5 trims since it causes known problems. …
@@
The Hyundai Kona has a very open and room-like interior, with some well-placed buttons. I’ve been on an ergonomics kick lately, closely scrutinizing the button placement and cockpit layouts of every car I get into. The 2024 Hyundai Kona has a uniquely open and homey-feeling interior, but one aspect of it that struck me was how perfect the heated seat and steering wheel buttons are. Before your eyes roll out of your head at the thought of praising something as mundane as a button, consider this: Heated seat and steering wheel controls are features you might use every day, multiple times a day, for months and months while you own a car. …
@@
Coverage: 4 Comprehensive Coverage of Main Issues: The insight summary effectively highlights the initial issues faced by the reviewer, namely the defective 12v battery and problematic software updates, and how these were resolved by upgrading to software version 2.1. … Inclusion of Recommendation and Software Concerns: The insight summary correctly encapsulates the reviewer’s advice against automatic updates due to the issues with version 2.2, emphasizing its buggy nature and incompatibility with B5 trims. … Minor Missing Details: Although the insight summary is generally accurate, it doesn’t mention the specific aspect of finding a stable software version and then turning off automatic updates, which adds a nuanced detail to the reviewer’s overall advice. …
@@
Specificity: 4 Resolution Details: The insight summary covers the resolution process involving the upgrade to software version 2.1 to address initial vehicle issues. … Specific Features and Versions: The insight summary specifies the difference between software versions 2.1 and 2.2, noting the inclusion of Apple CarPlay in version 2.2. … General Impression and Recommendations: While largely specific in detail, certain generalizations are made, such as mentioning the comfort and uniqueness of the SUV without referencing specific features that contribute to these impressions. …
@@
Insightfulness: 4 Comprehensive Coverage: The insight summary effectively covers the key points of the review, including the issue with the battery, the software problems, and the resolution through the 2.1 software update. … Value-Added Advice: By recommending a strategic approach to software updates and emphasizing the stability offered by version 2.1, the insight summary provides actionable insights that can directly inform decision-making for potential users of the XC60 B5, particularly those concerned about software reliability. Balanced Perspective: The insight summary maintains a balance by acknowledging the non-software-related positives of the vehicle, such as comfort and aesthetic appeal, which enriches understanding for readers interested in a broader evaluation of the SUV’s attributes. …
@@
Intent: 5 Tone and Intent Preservation: The insight summary effectively captures the reviewer’s tone of cautious appreciation towards the vehicle, highlighting the resolve to prioritize stability over newer, but potentially problematic, software updates. … Emotional Nuances: The emotional essence of the original review, which oscillates between initial frustration and eventual appreciation, is well-preserved. … Detail and Specificity: Key details concerning the software update issues, particularly the problems associated with version 2.2 and the decision to stick with version 2.1, are thoroughly covered in the insight summary. …
@@
Faithfulness: 5 Accurate Representation: The insight summary accurately captures the main points and perspectives shared by the commenters without distorting their meaning. … No Omitted Key Details: The core opinions about the use of ICE hulls for EVs, the practical considerations for work vans, and the limited design flexibility due to utility needs are all present in the insight summary. This shows no significant omissions. Absence of Hallucinated Information: The insight summary does not introduce any information that was not present in the original text, ensuring that it stays true to the opinions and insights provided by the commenters. …
@@
Specificity: 5 Concrete Alignment with Comments: The insight summary effectively captures the conversation between different users by directly referencing the points they make. It notes Person 1’s dissatisfaction with ICE-based EV designs and preference for innovative startups, Person 2’s emphasis on practicality and part availability for work vans, Person 3’s mention of financial and practical factors, and Person 4’s blunt assertion about vans’ inherent boxiness. … No Overgeneralization: At no point does the insight summary veer into vague or generic statements. Each opinion is not only linked to the respective commenter but also articulated with clear explanation of their rationale, demonstrating a detailed and precise understanding of the user opinions. Point-wise Clarity: The insight summary maintains a structurally coherent format by presenting each perspective distinctly. …
Insightfulness: 4 Comprehensive Overview: The insight summary effectively synthesizes various perspectives from the commentary, covering not only the dissatisfaction with traditional designs from Person 1 but also contrasting practical views from others who value utility over aesthetics. … Emphasis on Practical Factors: By highlighting the points about practicality—such as access to parts, financial logic, and the vans’ purpose—the insight summary goes beyond mere restatement of comments and offers a deeper understanding of the practical considerations influencing design decisions in the EV market. Concludes with Functional Insight: The insight summary concludes with Person 4’s point about the inherent boxy design of vans due to their function, driving home the idea that practicality often trumps aesthetic innovation in vehicle design. …
Intent: 5 Tone and Intent Preservation: The insight summary accurately captures the tone and intent of each comment within the discussion. Person 1’s dissatisfaction with the trend is clearly articulated, as is the practical-focused tone of Persons 2, 3, and 4. … Nuance and Specificity: The insight summary encapsulates the nuances in each participant’s perspective. It highlights Person 1’s preference for innovative design, contrasts this with the practical considerations highlighted by Person 2 and Person 3, and points out Person 4’s straightforward viewpoint regarding the inherent functionality of vans. Consistent Style and Balance: The insight summary maintains a balanced view and communicates both sides of the argument without straying from the original intent of the discourse. …